Changes between Initial Version and Version 1 of UC20200204


Ignore:
Timestamp:
Jul 30, 2021, 7:49:14 AM (10 months ago)
Author:
trac
Comment:

--

Legend:

Unmodified
Added
Removed
Modified
  • UC20200204

    v1 v1  
     1== APO 3.5-m Users Committee Telecon, 02/04/2020 ==
     2
     3Attending:  Nancy Chanover (NMSU), Mark Klaene (APO), Bill Ketzeback (APO), Kevin Schlaufman (JHU), Eric Bellm (UW), Ben Williams (UW), Zachory Berta-Thompson (CU), Misty Bentz (GSU), Moire Prescott (NMSU), Mukremin Kilic (OU), Joanne Hughes Clark (Seattle U), Russet !McMillan (APO), Kal Kadlec (UW), Chip Kobulnicky (UWy), Sarah Tuttle (UW)
     4
     5----
     6
     7=== User feedback and comments from institutional representatives: ===
     8
     9* NMSU - nothing to report
     10* Colorado - nothing to report
     11* Wyoming - nothing to report
     12* Oklahoma - nothing to report
     13* Washington - nothing to report
     14* Seattle - nothing to report
     15* FGCU - nothing to report (by email)
     16* Georgia State - nothing to report
     17* UVa - nothing to report (by email)
     18* JHU - nothing to report
     19* NAPG - no report
     20
     21----
     22
     23=== Telescope and Instruments Report: ===
     24
     25Mark's detailed report is included below, followed by additional information discussed during today's call.
     26
     27''3.5-m Telescope and Instruments Highlights, 11/29/2019 - 1/29/2020''
     28
     290) Overview
     30
     31Numerous winter storms have come through, some depositing several inches of snow at a time. However, a few have consisted primarily of rain so the moisture level overall has been good but snowpack is low. We did lose a little time due to snow/ice on the roof.  In a couple of instances there was no way to manually remove it, although this generally occurred during times of poor observing conditions.  Overall there have been no serious telescope or instrument issues. We continue to make substantial progress in planning for a public open house on March 7, with a half-night of time allocated for eyepiece viewing.
     32
     331) Telescope
     34
     35Nominal operations with minimal problems.
     36
     372) Instruments
     38
     39DIS: The blue camera contamination is still bad and there are no plans to service it. The red camera is at times worse and at times better. There hasn’t been a consistent worsening trend; we continue to monitor this.  There were minor issues with the cart and turret detent that were and are being addressed and monitored.
     40
     41!TripleSpec is operational. A pattern noise was noted at a low level. Ed found an issue with the connectors and that improved things quite a bit, although the noise level may still be present at times. We are back at the monitoring stage right now and trying to determine a repeatable cause.
     42
     43Agile is operational. There was a report of incorrect information in the header, which is being addressed.
     44
     45Echelle is operational. The inter-order light in the blue has significantly improved.
     46
     47NIC-FPS is operational. We are frequently getting readout errors that require an INIT to recover from. We recently reseated the cards in the controller and will monitor it to see if that helped.
     48
     49ARCTIC is operational. We had  to reset a lost quad. We also have a header issue that we are working on. We are working on improving the heater to make warmup safer and more efficient. The UW TEG is revisiting the issue of the diffuser rotation mechanism.
     50
     51ARCSAT: Has been operational. We are in the process of porting some computing tasks over to a new computer to hopefully reduce the frequency of computer crashes/hangs in the future.  We are also starting fabrication of a mount for the FGCU Shelyak echelle spectrograph. 
     52
     53----
     54
     55=== Additional telescope and instrument discussion: ===
     56
     57Mark provided a brief reminder of our winter operations mode. Because we have a flat enclosure roof, the day crew does go up for snow removal when conditions are safe enough to do so (safety considerations include ice, ice on the arrestors, and wind conditions). There are few but occasional times when we have to remain closed even when the sky conditions are such that we would like to be open because the enclosure roof could not be cleared.
     58
     59The telescope has been operating well. As has been discussed on previous calls, there are no plans to service DIS blue. The contamination in DIS red is not getting worse after all so we didn’t need to service it in January; we are still keeping an eye on it. There were recently some minor issues with the DIS turret and cart. The !TripleSpec noise is still intermittent and at a low level.  Work done on site may not have improved it; UC reps should warn their users that the noise pattern is there and they should let us know if it begins to impact their science. Russet has seen some correlations with rotator position so we need to know what rotation results in more noise; since it is intermittent the more examples we can get, the better.  The worst it has been is 20 DN peak to peak.  Users should check their data; it is aligned with the rows and columns of the chip. [Note added on 2/26/20: recent measurements during Engineering time suggest that the noise level is greatly reduced and now barely noticeable (7 DN peak to peak).  '''Users should still be made aware so they can let us know if it reoccurs.''']
     60
     61The IT staff has been working on header issues with Agile and ARCTIC. Nothing has changed, we're just trying to make them more accurate. In the case of ARCTIC, the exposure time reported in the header is always what was commanded and not the actual exposure time (e.g., if the chip is read out early, the exposure time does not update). The dark time is also not reported in the header, nor is chip temperature (the headers were recording the cold head temperature instead, which is not the chip temperature). For Agile, the headers are more complicated because part of it is written by the instrument and part is written by the hub (i.e., the software that ties the telescope control, instrument control, and TUI together). We knew that there were vulnerabilities in this shared header system when taking headers very rapidly (e.g. if the telescope is not tracking some of the header does not get properly written, image_type (OBJECT/BIAS/DARK) is always OBJECT since it does not have a shutter so there is no such thing as a true dark). Muk pointed out that as long as the timing information is accurate, users would probably be OK with some of these other issues.  Russet reminded us that the instrument has its own GPS clock so the timing should always be good.
     62
     63Regarding ARCSAT, Bill has been implementing some changes to share the computing load between two computers. He has also been working on the mount for the Shelyak spectrometer.  KS asked for some more info about the spectrograph.  It is an eShel (https://www.shelyak.com/produit/pf0011-eshel-spectroscope/?lang=en) spectrograph from Shelyak, on loan from  FGCU. We expect it will useful to the ARC community, both for student training and for science observations of bright targets.  The current plan is for it to be commissioned on the telescope in Q2, and we will provide a manual containing user instructions as well as anticipated performance specifications on the 0.5m.
     64
     65----
     66
     67=== Q1 3.5m scheduling: ===
     68
     69There is a little open time remaining late February and some in March. After today's schedule update there will be 5 time slots remaining (including some dark time).
     70
     71----
     72
     73=== Q1 ARCSAT scheduling: ===
     74
     75There is still lots of time remaining in February and March. We have received one request for some open time in February.  Contact Russet, Ben, and Nancy if interested in using it.
     76
     77----
     78
     79=== Q2 3.5m scheduling: ===
     80
     81The summer shutdown is expected to start on June 24, which will reduce the available time in Q2 by about a week.  We will be realuminizing the primary mirror this summer, and we are constrained by the availability of the realuminizing facility at Kitt Peak.   ZBT inquired about the throughput gain after realuminization.  Mark replied that it's about 2%, and there is also a significant reduction in scattered light.  Russet said that previously throughput measurements were made with DIS, but now they will do it with ARCTIC.  The obs-specs are working to get a good baseline now. We learned some lessons after realuminizing M2 and M3 two summers ago, namely that we need to capitalize on good weather during engineering time in the spring in order to complete these baseline measurements.
     82
     83----
     84
     85=== Q2 ARCSAT scheduling: ===
     86
     87The call for Q2 ARCSAT proposals will be issued once the Q2 3.5m schedule is completed.
     88
     89----
     90
     91=== NA2 Baffle Project: ===
     92
     93We now have the new NA2 baffle in hand. There were some issues after a fit test this morning so the site staff may need to make some minor modifications to it. The engineering tests will likely be done next week.  We have a commissioning plan, which could be executed in one night, weather permitting.
     94
     95----
     96
     97=== KOSMOS Update: ===
     98
     99The UW TEG is working on the dewar in lab to try to get it to hold vacuum.  They completed number of tests and it is now in the machine shop, where all of the O-ring seals are being ground to get the surface roughness down.  The detector is at e2V Teledyne in the UK, where it is being repaired. Estimated time for repair is two weeks. Conor is working on the controller software.
     100
     101----
     102
     103=== Status Update on Near-Term Action Items from Strategic Planning Meeting  ===
     104
     105* DAWG: The following update was provided by DAWG Chair Adam Kowalski:
     106
     107The APO Data Analysis Working Group (DAWG) was established at the strategic planning meeting in June 2019.  The group is composed of Jon Holtzman (NMSU), Jim Davenport (UW), Xinyu Dai (Univ. of Oklahoma), Sarah Tuttle (UW), Adam Kowalski (CU), Nancy Chanover, and Shane Thomas at APO.  We have met three times over Zoom between August 2019 - Jan 2020, and all meetings were recorded.  Our main task was to get feedback from ARC users on ways to improve the handling of 3.5m raw data, from archiving to post-processing (e.g., quick reduction tools and/or reduction guides for each instrument).
     108
     109We provided a Google survey to the ARC users in late 2019 in which we asked them to tell us how they reduced their data for each instrument and to rate their satisfaction with their data reduction method.   There were ~40 replies, and the results have been compiled and synthesized.  Overall, we found that there is still a significant base of IRAF/PyRAF users.  There is a general desire for each instrument page to provide at least a basic data reduction guide / recommendation, especially for ARCES.  Several users made us aware of promising avenues for IRAF/PyRAF replacements that could be used as reduction tools.
     110
     111The next step for the DAWG is to write a document that summarizes our findings from the users survey in which we point out the existence of such tools, which are not widely known yet.  We will provide this document to APO, in which we discuss the options, feasibility, and general user desire for quick-look / auto-reduction tools. We will also discuss the option for just providing reduction tool recommendations / links on each instrument webpage.  As the DAWG, we do not plan to make recommendations of specific tools/pipelines; instead, our main goal is to compile a list / inventory of all options to increase general awareness of what's out there.
     112
     113We also discussed the issue of data archiving and storage.  In our summary, we plan to include ideas for a proposal for a relatively cheap way to expand the storage of 3.5m data (to increase from 4 TB) and to extend the length of time over which data are stored for users within ARC. 
     114
     115* Getting narrow-band filters for ARCTIC: in progress. Nancy has selected a vendor.
     116
     117* Science communications interns: The two student interns are increasing the social media presence of APO. They are starting by writing short summaries of some recent papers (accepted or published) that make use of 3.5m data. They will also be doing a staff highlights segment where they introduce various staff members to our followers on social media.  If there is anything you would like to see featured on the APO Facebook page or Twitter account, please let Nancy know!
     118
     119* AAS recap: Nancy reported that the first 3.5m exhibit booth was a success. A number of people stopped by and inquired about the capabilities of the 3.5m (instruments, scheduling, etc.). This was an important first step in getting the word out about APO, and we will likely continue doing this in the future (winter AAS meetings only).  Special thanks to everyone who volunteered some time at the booth and spoke with interested parties about APO.
     120
     121----
     122
     123=== ACTION ITEMS:===
     124
     125'''Open action items from previous meetings:'''
     126
     127* ''Nancy will contact trip leaders bringing student groups to APO in Q4 (-> Q1) to discuss training video ideas.''  STATUS: OPEN.
     128
     129'''New action items from this meeting:'''
     130
     131None.
     132
     133----
     134
     135=== All Other Business ===
     136
     137KS asked who would be the appropriate POC for arranging a site tour for some JHU students who would be visiting the southwest this spring. Mark said he could be that person.
     138
     139Mark announced that we will be holding a community Open House event on March 7, which will include daytime outreach activities at APO, evening lightning talks at the Sunspot Astronomy and Visitors Center, and nighttime tours of the 3.5 (including eyepiece viewing!) on the hour between 6 pm and 12 am.  Nearly all tour slots are already booked, and planning is going well.
     140
     141----
     142
     143=== Next meeting:===
     144
     145The next Users Committee Zoom meeting will be on Monday Mar. 2 at 10:30 am MST.
     146
     147
     148----